# Search for exotic physics at BESIII

### v1

@@ 2.2 Search for the massless dark photon
which includes the NP energy scale $\Lambda_{\rm{NP}}$ and dimensionless coefficients $C_{jk}$. In this operator, it has a naturally flavor-violating coupling, such as $cu\gamma^{\prime}$ coupling. Based on $4.5\leavevmode\nobreak\ \rm{fb}^{-1}$ $e^{+}e^{-}$ annihilation data at $\sqrt{s}=4.6\sim 4.7$ GeV, the first search for the $cu\gamma^{\prime}$ coupling is performed through the process $\Lambda_{c}\to p\gamma^{\prime}$, where the invisible massless dark photon $\gamma^{\prime}$ is detected from the recoiling. No significant signals are observed, and the UL of the branching fraction (BF) is set to be $8.0\times 10^{-5}$ at the 90% CL [12], corresponding to the NP energy scale related parameter $|\mathcal{C}|^{2}+|\mathcal{C}_{5}|^{2}<9.6\times 10^{-16}\rm{GeV}^{-2}$, where $\mathcal{C}=\Lambda_{\rm{NP}}^{-2}(C^{u}_{12}+C^{u*}{12})\nu/\sqrt{8}$, $\mathcal{C}_{5}=\Lambda_{\rm{NP}}^{-2}(C^{u}_{12}-C^{u*}{12})\nu/\sqrt{8}$, and $\nu$ is Higgs vacuum expected value.
@@ 2.7 Recent glueball studies
$J/\psi$ radiative decays provide gluon rich environment and an ideal place for glueball study. A partial wave analysis (PWA) of $J/\psi\to\gamma K_{S}K_{S}\pi^{0}$ is performed at BESIII, and two pseudo-scalar states are observed, $\eta(1405)$ and $\eta(1475)$ [21], as shown in Fig 5 (a). In the Quark model prediction, there is only one pseudo-scalar meson near 1.4 GeV, and $\eta(1475)$ is regarded as the first radial excitation of $\eta^{\prime}$ while $\eta(1405)$ may be a pseudo-scalar glueball candidate. This measurement could be an important input for $0^{-+}$ glueball. In the LQCD prediction, the mass of $0^{-+}$ glueball is in the range of 2.3 $\sim$ 2.4 $\rm{GeV}/c^{2}$, and some theorists attempt to explain $\eta(1405)$ and $\eta(1475)$ using one pole [20]. Further studies are needed.

### v2

@@ 2.2 Search for the massless dark photon
which includes the NP energy scale $\Lambda_{\rm{NP}}$ and dimensionless coefficients $C_{jk}$. In this operator, it has a naturally flavor-violating coupling, such as $cu\gamma^{\prime}$ coupling. Based on $4.5\leavevmode\nobreak\ \rm{fb}^{-1}$ $e^{+}e^{-}$ annihilation data at $\sqrt{s}=4.6\sim 4.7$ GeV, the first search for the $cu\gamma^{\prime}$ coupling is performed through the process $\Lambda_{c}\to p\gamma^{\prime}$, where the invisible massless dark photon $\gamma^{\prime}$ is detected from the recoiling. No significant signals are observed, and the UL of the branching fraction (BF) is set to be $8.0\times 10^{-5}$ at the 90% CL [12], corresponding to the NP energy scale related parameter $|\mathcal{C}|^{2}+|\mathcal{C}_{5}|^{2}<9.6\times 10^{-16}\rm{GeV}^{-2}$, where $\mathcal{C}=\Lambda_{\rm{NP}}^{-2}(C^{u}_{12}+C^{u*}{12})\nu/\sqrt{8}$, $\mathcal{C}_{5}=\Lambda_{\rm{NP}}^{-2}(C^{u}_{12}-C^{u*}{12})\nu/\sqrt{8}$, and $\nu$ is Higgs vacuum expected value [13].
@@ 2.7 Recent glueball studies
$J/\psi$ radiative decays provide gluon rich environment and an ideal place for glueball study. A partial wave analysis (PWA) of $J/\psi\to\gamma K_{S}K_{S}\pi^{0}$ is performed at BESIII, and two pseudo-scalar states are observed, $\eta(1405)$ and $\eta(1475)$ [21], as shown in Fig 5 (a). In the Quark model prediction, there is only one pseudo-scalar meson near 1.4 GeV, and $\eta(1475)$ is regarded as the first radial excitation of $\eta^{\prime}$ while $\eta(1405)$ may be a pseudo-scalar glueball candidate. This measurement could be an important input for $0^{-+}$ glueball. In the LQCD prediction, the mass of $0^{-+}$ glueball is in the range of 2.3 $\sim$ 2.4 $\rm{GeV}/c^{2}$. Some theorists attempt to explain $\eta(1405)$/$\eta(1475)$ using a single state [22] , while others propose more than one pole [23].